News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 9 March 2001 10.00am EST

During the last week, Chandra completed the observing schedule as planned.

Chandra continued through the 2001 Spring eclipse season with nominal power and thermal performance. A real-time procedure was completed on March 8 to uplink the software patch in preparation for the 11th and final (short) eclipse of the season that will occur this afternoon.

Of note next week is the 15 March deadline for the submission of proposals for Chandra's third observing cycle. The Peer Review selection is scheduled for June and observations from this cycle are expected to begin in November 2001.

The planned observations for the coming week are shown below and include an observation of HD 111812 coordinated with the Hubble Space Telescope. The coordinated portion of the observation will occur on Mar 13 between 8:23am and 2:04pm EST.

------------------------------------------
HD 111812               ACIS-S/HETG  Mar 12
KI 284                  ACIS-S       Mar 14
Crab Nebula             ACIS-S
Radiation Belt
NGC 4507                ACIS-S/HETG  Mar 15
SS 433                  ACIS-S/HETG  Mar 16
Radiation Belt                       Mar 17
NGC 2865                ACIS-S
RCW103                  ACIS-I       Mar 18
G21.5-09, S3, AO2A      ACIS-S
LBQS 1331-0108          ACIS-S
SDSS120823+001027       ACIS-S
LBQS 1235+0857          ACIS-S
LBQS 1216+1103          ACIS-S

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports